$SPAD/src/input cone.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- graph of a cone
draw(surface(u*cos(v),u*sin(v),u),u = -3..3,v = 0..2*%pi,[title "Cone"])

———-

References^{[1] nothing}

4